31-5 In an oscillating LC circuit, L = 1.64 mH and C = 3.91 ?F. The...

31-5

In an oscillating LC circuit, L = 1.64 mH and C = 3.91 ?F. The maximum charge on the capacitor is 3.09 ?C. Find the maximum current.
Answer #1

Here ,

L = 1.64 mH

C = 3.91 uF

Q = 3.09 uC

let the maximum current is I

0.50 * L * I^2 = 0.50 * Q^2/C

0.50 * 1.64 *10^-3 * I^2 = 0.50 * (3.09 * 10^-6)^2/(3.91 *10^-6)

solving for I

I = 0.039 A

the maximum current is 0.039 A